# Conditional-$t^{3}$VAE: Equitable Latent Space Allocation for Fair Generation

### v1

@@ 5.2.3 Per-class evaluation
Figure 5: Fine-grained comparison of C-$t^{3}$VAE and C-VAE models under varying imbalance ratios on SVHN-LT.
On the CelebA dataset (Figure 4), C-$t^{3}$VAE improves Recall and F1 on the most imbalanced attribute (Mustache), but not on more balanced ones (Male, Smiling) which is an expected behavior. However, surprisingly, for the attribute Young (imbalance ratio 3.5), C-VAE performs slightly better this suggests that there is a regime where Gaussian priors may still suffice. To explore this, we vary the imbalance ratio on SVHN-LT from $100$ to $1$ and plot the results in Figure 5. This figure shows there is a threshold $\rho\approx 3$ before which C-VAE performs better; beyond that, C-$t^{3}$VAE has the advantage. Also, we notice that the more the imbalance ratio increases, the more FID gap becomes more significant and in favor of the C-$t^{3}$VAE.
For the SVHN-LT dataset (Figures in Appendix G), when comparing the C-$t^{3}$VAE to the C-VAE, the C-$t^{3}$VAE achieves significantly higher Recall and competitive Precision. This indicates a better coverage of data modes and maintaining images quality which yields higher F1 scores especially on tail classes. For the CIFAR100-LT dataset (Figures in Appendix G), C-VAE often achieves high Precision but zero Recall, indicating mode collapse. In contrast, the C-$t^{3}$VAE preserves Recall at the cost of slightly lower Precision, also resulting in improved F1 scores. Consequently, on the SVHN-LT CIFAR100-LT and CelebA datasets, C-$t^{3}$VAE sets itself as the most reliable method for fair, high quality image generation.
@@ 6 Conclusion
We introduced C-$t^{3}$VAE, a class-conditional generative model that leverages Student’s t-distributions in the latent space with a theoretically derived balanced sampling scheme. This approach improves fairness and sample quality under class imbalance. Extensive experiments on SVHN-LT, CIFAR100-LT, and CelebA demonstrate that, after the optimization of the $\beta$, $\nu$ and $\tau$ hyper-parameters, C-$t^{3}$VAE consistently outperforms the $t^{3}$VAE and conditional VAE baselines, achieving up to a 15-point FID improvement on highly imbalanced datasets. Per-class Precision, Recall, and F1 evaluations confirm improved mode coverage, particularly in tail classes. Moreover, we identify $\rho>3$ as the regime where C-$t^{3}$VAE outperforms Gaussian-prior models, with the performance gap widening as $\rho$ increases. For $\rho\lesssim 3$, C-$t^{3}$VAE remains competitive with C-VAE.
Beyond generative performance, C-$t^{3}$VAE contributes to fairness-aware modeling by promoting balanced latent sampling and better mode coverage in underrepresented classes. Future work will explore extending C-$t^{3}$VAE to multi-label problems and integration with Latent Diffusion models.
Acknowledgement
The authors were granted access to the HPC resources of MesoPSL financed by the Région Île-de-France and the Equip@Meso project (reference ANR-10-EQPX-29-01) of the programme investissements d’avenir supervised by France’s Agence nationale pour la recherche.
Appendix for
”Conditional-$t^{3}$VAE: Equitable Latent Space Allocation for Fair Generation”
Appendix A $\gamma$-power divergence corrected derivation
In this section, we present our correction of the previously derived form of the $\gamma$-power divergence in [10]. We start from the integrals derived in [10] and keep a similar notation :
$\int q(x)p(x)^{\gamma}dx$ $=C_{\nu,d}^{\gamma}|\Sigma_{1}|^{-\frac{\gamma}{2}}\left(1+\frac{1}{\nu-2}\operatorname{Tr}(\Sigma_{1}^{-1}\Sigma_{0})+\frac{1}{\nu}(\mu_{0}-\mu_{1})^{\top}\Sigma_{1}^{-1}(\mu_{0}-\mu_{1})\right)$
$\int q(x)^{1+\gamma}dx$ $=C_{\nu,d}^{\gamma}|\Sigma_{0}|^{-\frac{\gamma}{2}}\left(1+\frac{d}{\nu-2}\right)$
$\int p(x)^{1+\gamma}dx$ $=C_{\nu,d}^{\gamma}|\Sigma_{1}|^{-\frac{\gamma}{2}}\left(1+\frac{d}{\nu-2}\right)$
Combining these formulas gives a similar enthropy $\mathcal{H}_{\gamma}$ to the one derived in [10]:
$\mathcal{H}_{\gamma}(q)=-C_{\nu,d}^{\frac{\gamma}{1+\gamma}}|\Sigma_{0}|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{d}{\nu-2}\right)^{\frac{1}{1+\gamma}}.$
However, the cross-entropy $\mathcal{C}_{\gamma}$ takes the following form :
[…]
In red, we highlight the main difference to the formula derived in [10]. $\mathcal{H}_{\gamma}(q)$ and $\mathcal{C}_{\gamma}(q,p)$ combine to give :
$\mathcal{D}_{\gamma}(q\|p)$ $=-\frac{C_{\nu,d}^{\frac{\gamma}{1+\gamma}}}{\gamma}\left(1+\frac{d}{\nu-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg{[}-|\Sigma_{0}|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{d}{\nu-2}\right)$
[…]
Appendix B Priors derivations
@@
$=\int C_{\nu,m+n}|\Sigma_{x}|^{-\frac{1}{2}}|\Sigma_{y}|^{-\frac{1}{2}}\Bigg{[}1+\frac{(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})}{\nu}$
$\quad+\frac{(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\nu}\Bigg{]}^{-\frac{\nu+m+n}{2}}dx$
@@
$\times\int\left(1+\frac{(1+\nu^{-1}m)(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)(\nu+m)}\right)^{-\frac{\nu+m+n}{2}}dx.$
@@
Appendix C Loss function derivation
@@
$\quad+\frac{(\mu(x)-\mu_{y})^{\top}\Sigma_{y}^{-1}(\mu(x)-\mu_{y})}{\nu}$
$\quad+\frac{\mathbb{E}_{z}\left[(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))\right]}{\nu}\Bigg{]}.$
@@
Appendix D Sampling distribution variance derivation
@@
$\mathcal{D}(q\|p^{\star})$ $=-\frac{C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}}{\gamma}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg{[}-\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}$
$\quad\times\left(1+\frac{m}{\nu+n-2}\right)+|\tau^{2}I|^{-\frac{\gamma}{2}}|\left(1+\nu^{-1}n\right)^{-1}\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}$
$\quad\times\left(1+\frac{\operatorname{Tr}\left(\tau^{-2}\left(1+\nu^{-1}n\right)^{-1}\Sigma_{\phi}(x)\right)}{\nu+n-2}+\frac{\tau^{-2}}{\nu+n}\|\mu(x)-\mu_{y})\|^{2}\right)\Bigg{]}$
@@
$\mathcal{D}(q\|p^{\star})$ $=-\frac{1}{\gamma}C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg{[}-\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}$
$\quad\times\left(1+\frac{m}{\nu+n-2}\right)+\tau^{-\gamma m}\left(1+\nu^{-1}n\right)^{-\frac{\gamma^{2}m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}$
$\quad\left(1+\frac{\tau^{-2}\left(1+\nu^{-1}n\right)^{-1}}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)+\frac{\tau^{-2}}{\nu+n}\|\mu(x)-\mu_{y})\|^{2}\right)\Bigg{]}$
$=-\frac{1}{\gamma}C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg{[}-\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}$
$\quad\times\left(1+\frac{m}{\nu+n-2}\right)+\nu^{-1}\tau^{-2-\gamma m}\left(1+\nu^{-1}n\right)^{-\frac{\gamma^{2}m}{2(1+\gamma)}-1}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}$
$\quad\left(\kappa+\frac{\nu}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)+\|\mu(x)-\mu_{y})\|^{2}\right)\Bigg{]}$
$=-\frac{1}{\gamma}C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\nu^{-1}\tau^{-2-\gamma m}\left(1+\nu^{-1}n\right)^{-\frac{\gamma^{2}m}{2(1+\gamma)}-1}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}$
$\quad\times\Bigg{[}-\nu\tau^{2+\gamma m}\left(1+\nu^{-1}n\right)^{\frac{\gamma^{2}m+\gamma m}{2(1+\gamma)}+1}|\Sigma_{\phi}(x)|^{-\frac{\gamma^{2}}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)$
$\quad+\kappa+\frac{\nu}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)+\|\mu(x)-\mu_{y})\|^{2}\Bigg{]},$
@@
$\tau^{2+\gamma m}\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2}+1}|\Sigma_{\phi}(x)|^{-\frac{\gamma^{2}}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)=\frac{C_{1}}{C_{2}}.$
@@
$\tau^{2+\gamma m}$ $\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2}+1}|\Sigma_{\phi}(x)|^{-\frac{\gamma^{2}}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\;\,=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\frac{m}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
$\tau^{2+\gamma m}$ $=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\nu^{-1}n\right)^{\frac{-\gamma m}{2}-1}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)^{-1}\left(1+\frac{m}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}$
$\quad\times\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
$=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\nu^{-1}n\right)^{-\frac{\gamma m}{2}-1}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
$=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\nu^{-1}n\right)^{-\frac{\gamma m}{2}-1}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}\left(\frac{\nu+n-2}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
$=\left(1+\nu^{-1}n\right)^{-\frac{\gamma m}{2}-1}\left(\sigma^{n}C^{-1}_{\nu,n}|\Sigma_{\phi}(x)|^{\frac{\gamma}{2}}\quad\frac{\nu+n-2}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}.$
@@
$\tau^{2}$ $=\left(1+\nu^{-1}n\right)^{-\frac{\gamma m+2}{\gamma m+2}}\left(\sigma^{n}C^{-1}_{\nu,n}|\Sigma_{\phi}(x)|^{\frac{\gamma}{2}}\frac{\nu+n-2}{\nu-2}\right)^{\frac{2\gamma}{(1+\gamma)(2+\gamma m)}}$
$=\left(1+\nu^{-1}n\right)^{-1}\left(\sigma^{n}C^{-1}_{\nu,n}|\Sigma_{\phi}(x)|^{\frac{\gamma}{2}}\frac{\nu+n-2}{\nu-2}\right)^{\frac{2\gamma}{(1+\gamma)(2+\gamma m)}},$
@@
Appendix E Experimental Setup Details
E.1 Datasets
We conduct experiments on three datasets notably SVHN-LT [11], CIFAR100-LT [24, 12] and CelebA [13] each chosen to highlight different challenges related to generative modeling under class imbalance and varying visual complexity.
  • SVHN-LT : This dataset is comprised of colored images of digits from 0 to 9 of size $32\times 32\times 3$. It serves as our controlled experimental setting. While simple enough for all models to converge, it is rich enough to reflect performance differences. However, as this dataset is naturally imbalanced, we balance the number of images across classes to have more control over the imposed imbalance ratio. In Table 2 we provide the number of images present in the dataset before balancing.

### v2

@@ 5.5 Per-class evaluation
On CelebA (Figure 4), C-$t^{3}$VAE improves Recall and F1 on the most imbalanced attribute (Mustache), but not on more balanced ones (Male, Smiling), which is expected. When class imbalance is moderate (e.g., Young attribute with $\rho=3.5$), C-VAE slightly outperforms C-$t^{3}$VAE, indicating that Gaussian priors suffice when encoder variance and decoder flexibility can compensate for small frequency skew. This suggests the presence of a regime where Gaussian priors may suffice. To explore this, we vary the imbalance ratio on SVHN-LT from $100$ to $1$ and plot the results in Figure 5. This figure shows a threshold around $\rho\approx 5$: below it C-VAE can be better, while above it C-$t^{3}$VAE has the advantage. As the imbalance ratio increases, the FID gap grows in favor of C-$t^{3}$VAE. For moderate imbalance ($\rho<5$), encoder variance and decoder flexibility can compensate for frequency skew; beyond this regime, posterior–prior mismatch grows nonlinearly, and Gaussian priors increasingly compress minority-class latent regions. Heavy-tailed class-conditional priors mitigate this compression by permitting larger dispersion without penalizing minority likelihood, explaining the observed transition.
Figure 5: C-$t^{3}$VAE versus C-VAE models under varying imbalance ratios on the SVHN-LT dataset.
For SVHN-LT (Figures in Appendix F), C-$t^{3}$VAE achieves higher Recall and competitive Precision compared to C-VAE, indicating better mode coverage while maintaining image quality, which yields higher F1 scores, especially on tail classes. For CIFAR100-LT (Figures in Appendix F), C-VAE often attains high Precision but near-zero Recall, indicating mode collapse. In contrast, C-$t^{3}$VAE preserves Recall at the cost of slightly lower Precision, resulting in improved F1 scores. Consequently, on SVHN-LT, CIFAR100-LT, and CelebA, C-$t^{3}$VAE is the most reliable method in our study for high-imbalance settings within the VAE family.
@@ 6 Conclusion
In this work, we introduced C-$t^{3}$VAE. This class-conditional generative model mitigates frequency-aligned latent allocation effects observed under imbalance. It uses per-class Student’s $t$-distributions in the latent space, paired with a theoretically derived, equal-weight sampling scheme. This design allocates uniform prior mass across all classes while capturing heavy-tailed intra-class variations. Furthermore, these structural improvements do not sacrifice efficiency. The computational complexity of C-$t^{3}$VAE matches that of standard conditional VAEs.
We evaluated our model on SVHN-LT, CIFAR100-LT, and CelebA. After optimizing the hyperparameters $\beta$, $\nu$, and $\tau$, C-$t^{3}$VAE consistently outperforms baseline models in high-imbalance regimes. It achieves up to a 15-point FID improvement over $t^{3}$VAE and conditional VAEs. Additionally, per-class Precision, Recall, and F1 metrics confirm superior mode coverage for tail classes. We also identified a critical imbalance ratio threshold at $\rho\approx 5$. For milder imbalances ($\rho\lesssim 5$), Gaussian-based models remain competitive. However, as imbalance grows beyond this threshold, C-$t^{3}$VAE demonstrates a significant and widening advantage.
Future research should explore extending the C-$t^{3}$VAE framework to multi-label settings. The MultiFacet VAE model (Falck et al., 2021) provides a promising starting point for handling complex, overlapping annotations. Furthermore, our experiments show that the sampling variance $\tau$ heavily impacts image quality. This parameter currently requires dataset-specific tuning. Therefore, developing tighter theoretical bounds and adaptive latent-space sampling methods remains a crucial next step for high-fidelity VAE generation.
Acknowledgment
This work received access to the High-Performance Computing (HPC) resources of MesoPSL, financed by the Region Île-de-France and the Equip@Meso project (reference ANR-10-EQPX-29-01) of the Investissements d’avenir program supervised by the Agence nationale pour la recherche.
Supplementary Material
Appendix A Priors derivations
@@
$=\int C_{\nu,m+n}|\Sigma_{x}|^{-\frac{1}{2}}|\Sigma_{y}|^{-\frac{1}{2}}\Bigg[1+\frac{(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})}{\nu}+\frac{(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\nu}\Bigg]^{-\frac{\nu+m+n}{2}}dx$
@@
$\quad\times\int\left(1+\frac{(1+\nu^{-1}m)(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)(\nu+m)}\right)^{-\frac{\nu+m+n}{2}}dx.$
@@
Appendix B Loss function derivation
@@
$\quad+\frac{(\mu(x)-\mu_{y})^{\top}\Sigma_{y}^{-1}(\mu(x)-\mu_{y})}{\nu}+\frac{\mathbb{E}_{z}\left[(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))\right]}{\nu}\Bigg].$
@@
Appendix C Sampling distribution variance derivation
@@
$\mathcal{D}(q\|p^{\star})$ $=-\frac{C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}}{\gamma}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg[-\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)$
$\quad+|\tau^{2}I|^{-\frac{\gamma}{2(1+\gamma)}}\times\left(1+\frac{\operatorname{Tr}\left(\tau^{-2}\left(1+\nu^{-1}n\right)^{-1}\Sigma_{\phi}(x)\right)}{\nu+n-2}+\frac{\tau^{-2}}{\nu+n}\|\mu(x)-\mu_{y})\|^{2}\right)\Bigg]$
@@
$\mathcal{D}(q\|p^{\star})$ $=-\frac{1}{\gamma}C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg[-\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)$
$\quad+\tau^{\frac{-\gamma m}{1+\gamma}}\left(1+\frac{\tau^{-2}\left(1+\nu^{-1}n\right)^{-1}}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)+\frac{\tau^{-2}}{\nu+n}\|\mu(x)-\mu_{y})\|^{2}\right)\Bigg]$
$=-\frac{1}{\gamma}C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg[-\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)$
$\quad+\frac{1}{\nu+n}\tau^{-2-\frac{\gamma m}{1+\gamma}}\left(\kappa+\frac{\nu}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)+\|\mu(x)-\mu_{y})\|^{2}\right)\Bigg]$
$=-\frac{1}{\gamma}C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\frac{1}{\nu+n}\tau^{-2-\frac{\gamma m}{1+\gamma}}\Bigg[-(\nu+n)\tau^{2+\frac{\gamma m}{1+\gamma}}\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}$
$\quad|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)+\kappa+\frac{\nu}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)+\|\mu(x)-\mu_{y})\|^{2}\Bigg],$
@@
$\tau^{2+\frac{\gamma m}{1+\gamma}}\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}+1}\left(1+\frac{m}{\nu+n-2}\right)=\frac{C_{1}}{C_{2}}.$
@@
$\tau^{2+\frac{\gamma m}{1+\gamma}}\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}+1}\left(1+\frac{m}{\nu+n-2}\right)=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\frac{m}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
$\tau^{2+\frac{\gamma m}{1+\gamma}}$ $=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\nu^{-1}n\right)^{-\frac{\gamma m}{2(1+\gamma)}-1}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
$=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\nu^{-1}n\right)^{-\frac{\gamma m}{2(1+\gamma)}-1}\left(\frac{\nu+n-2}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
$=\left(1+\nu^{-1}n\right)^{-\frac{\gamma m}{2(1+\gamma)}-1}\left(\sigma^{n}C^{-1}_{\nu,n}\quad\frac{\nu+n-2}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}.$
@@
$\tau^{2}=\left(1+\nu^{-1}n\right)^{-1}\left(\sigma^{n}C^{-1}_{\nu,n}\quad\frac{\nu+n-2}{\nu-2}\right)^{-\frac{2}{\nu+n-2}}.$
@@
Appendix D Experimental setup
D.1 Datasets
We conduct experiments on three datasets notably SVHN-LT (Netzer et al., 2011), CIFAR100-LT (Krizhevsky, 2009; Cao et al., 2019) and CelebA (Liu et al., 2015) each chosen to highlight different challenges related to generative modeling under class imbalance and varying visual complexity.
  • SVHN-LT : The Street View House Numbers (SVHN) dataset (Netzer et al., 2011) is composed of real-world digit images collected from Google Street View. It contains more than 600,000600{,}000 labeled digits (099) with size $32\times 32$, complex backgrounds, and diverse illumination conditions. The digits in SVHN are not centered or uniformly scaled, which makes the dataset considerably more challenging. Figure 6 shows sample images of this dataset.
    Figure 6: Sample images from the SVHN dataset (Netzer et al., 2011).
    However, as this dataset is naturally imbalanced, we balance the number of images across classes to have full control over the imposed imbalance ratio. In Table 2 we provide the number of images present in the dataset before balancing.